Suppose that x liters of pure peroxide are added to 700 liters of a 20% peroxide solution. Write a formula that gives the concentration, C, of the new mixture. How many liters of pure peroxide should be added to produce a new mixture that is 35% peroxide? Round to the nearest liter. Please show work.
Answer by richwmiller(17219)

.20*700+x=c*(700+x)
.20*700+x=.35*(700+x)
x = 161.538
x=162 rounded